### v1

@@ Abstract
Style transfer is a promising approach to close the sim-to-real gap in medical endoscopy. Rendering realistic endoscopic videos by traversing pre-operative scans (such as MRI or CT) can generate realistic simulations as well as ground truth camera poses and depth maps. Although image-to-image (I2I) translation models such as CycleGAN perform well, they are unsuitable for video-to-video synthesis due to the lack of temporal consistency, resulting in artifacts between frames. We propose MeshBrush, a neural mesh stylization method to synthesize temporally consistent videos with differentiable rendering. MeshBrush uses the underlying geometry of patient imaging data while leveraging existing I2I methods. With learned per-vertex textures, the stylized mesh guarantees consistency while producing high-fidelity outputs. We demonstrate that mesh stylization is a promising approach for creating realistic simulations for downstream tasks such as training and preoperative planning. Although our method is tested and designed for ureteroscopy, its components are transferable to general endoscopic and laparoscopic procedures.
@@
Mesh Stylization Endoscopy Differentiable Rendering Consistent Style Transfer.
@@ 1 Introduction
Figure 1: Sample camera trajectory inside the mesh. From top to bottom: original rendered, I2I style transfers [19], MeshBrush. Note inconsistency from frame to frame in the I2I method while MeshBrush maintains consistent features.
@@
To the best of our knowledge, our work is the first in endoscopic mesh stylization to synthesize consistently style-transferred videos from rendered data. Because mesh textures are stationary, MeshBrush guarantees consistency regardless of time and view. Leveraging existing I2I style transfer modules, our self-supervised network simply requires an input mesh of the patient anatomy and a pretrained style transfer module. Our lightweight model produces high-resolution spatial textures on the patient mesh to create realistic endoscopic sequences with temporal and global consistency using a view-dependent heatmap loss. Finally, we demonstrate that temporally consistent style transfer is effective in vision algorithms such as feature matching and Structure-from-Motion.
@@ 2 Related Work
Few works address temporal and long-term consistent style transfer in medical scenes. Frisch et al. use optical flow (RAFT [27]) to warp views and impose a time-invariant structural similarity loss in cataract surgical video [8]. However, as flow estimates are imperfect, the results are suboptimal. Engelhardt et al. extend CycleGAN [31] to process multiple consecutive frames simultaneously for multi-frame consistency in phantom surgical videos [6], which cannot enforce long-term consistency, e.g. when the endoscope returns to a location after a certain duration. Rivoir et al. use the underlying 3D model to impose similar textures via a view consistency loss and learnable neural textures to ensure global consistency in simulated laparoscopy [21]. Although this method demonstrates remarkable performance, it does not guarantee realistic and consistent translated views outside its training set. Mesh stylization, conversely, ensures consistency regardless of time, viewpoint, and novel views.
@@
3 Methods
Figure 2 shows the overview of MeshBrush. First, a mesh is segmented from a CT scan with medical visualization toolkits such as 3D-Slicer [^1] https://www.slicer.org/. We then generate a mesh skeleton via [2], from which $N$ camera views are sampled with the farthest point down sampling. The camera view direction is set to be along its line segment of the skeleton.
We increase the number of vertices of the input mesh by subdividing each triangle, increasing the resolution of the learned textures. From each camera pose $T_{i}\in SE(3)$, a differentiable rendering module renders a synthetic, unrealistic image $A_{i}^{(0)}$ at timestep 00. This image undergoes style transfer from an I2I pretrained model [19] to generate a realistic endoscopic view, $B_{i}$. Our modularized approach allows our framework to be used for various medical domains as users can change the I2I style transfer method to match each domain. The style transferred image $B_{i}$, generated from $A_{i}^{(0)}$, supervises the mesh textures of the vertices in view from $T_{i}$. At each iteration, this is repeated for each camera pose $T_{i}$, $i=1,...,N$.
@@ 3 Methods
During training, the vertex coordinates $V\in\mathbb{R}^{N_{v}\times 3}$ of the mesh are passed into a multilayer perceptron (MLP) to regress RGB values $C\in\mathbb{R}^{N_{v}\times 3}$ for each vertex. As MLPs are not proficient in learning high-frequency associations such as texture variations, we encode our input vertex coordinates to a Fourier feature representation to encourage the model to learn finer details of our mesh [26], i.e. $\gamma(V)=(\cos(2\pi\textbf{B}V),\sin(2\pi\textbf{B}V))$, where $B\in\mathbb{R}^{N_{v}\times 3}$ is a matrix with elements drawn from a Gaussian distribution. The texture predictions from the MLP are applied to the mesh, and new views are rendered from $T_{i}$ to generate $A_{i}^{(1)}$ from inside the anatomy with imperfect textures. Via differentiable rendering, the MLP learns valid vertex textures by backpropagating the loss between $A_{i}^{(1)}$ and $B_{i}$, expounded in the next section. This continues over many iterations until the error between $A_{i}^{(t)}$ and $B_{i}$ is minimized through training, which causes the mesh textures to mimic the style-transferred images.
@@
Using a neural network to predict textures rather than projecting the style transfer RGB values to the mesh has important advantages. First, vertices are much spatially finer than pixels, which enables the mesh to learn more realistic textures at a higher resolution. In addition, projecting RGB values from $B_{i}$ introduces quilting, where camera poses with overlapping image regions introduce sharp texture discontinuities due to inconsistency in the transferred styles.
@@
Given a rendered image $A\in\mathbb{R}^{H\times W}$, we first calculate $P\in\mathbb{R}^{H\times W\times 3}$, where $P_{i,j}$ represents the normal of the vertex closest to the camera at pixel $(i,j)$. We calculate the dot product between $P_{i,j}$ and the camera viewing direction $T_{f}\in\mathbb{R}^{3}$ for each pixel in the image to encode the angle $\theta$ between the vertex normals and the viewing direction. This creates the view-orientation heatmap $\Theta\in\mathbb{R}^{H\times W}$ with ranges of values from $\left[-1,1\right]$. Similarly, the view-distance heatmap between the camera and closest vertex at a given pixel $(i,j)$ is $\mathbb{D}=|X_{i,j}-T_{c}|_{2}$ where $X_{i,j}$ is the coordinate of the vertex closest to the camera at pixel $(i,j)$, and $T_{c}$ is the camera center in world coordinates. After scaling $\mathbb{D}$ to be in the range of $\left[0,1\right]$, the heatmaps are combined to create the aggregated heatmap $H\in\left[0,1\right]$.
$H=\frac{1}{2}\left(\frac{1-\Theta}{2}+\mathbb{D}\right)$ (1)
The final heatmap is generated by forcing pixels with view-direction values greater than a user-defined threshold to 0, and $A_{i}^{t}$ and $B_{i}$ are converted into LAB color space for more accurate color predictions. This results in our final heatmap-weighed MSE loss function:
$\mathcal{L}=H\cdot|A_{i}^{(t)}-B_{i}|^{2}_{2}$ (2)
@@ 5 Discussion and Conclusion
Structure-from-Motion. The sparse reconstructions of COLMAP, for both PyTorch3D and Pyrender renderings of the mesh, failed for both untextured and I2I methods. This confirms that textures from untextured and I2I-translated images do not provide unique descriptors for global camera localization. On the other hand, the features predicted by MeshBrush have globally distinct descriptors. We found that COLMAP uses 41.3% from 1,120 images to reconstruct 32327 points. Previous work achieved an average of 20.6% from 498 images to reconstruct 4838 points for kidney reconstructions from rendered views [1]. Thus, our method ensures the global and long-term consistency necessary for sophisticated downstream vision tasks like 3D reconstruction.

### v2

@@ Abstract
Style transfer is a promising approach to close the sim-to-real gap in medical endoscopy. Rendering synthetic endoscopic videos by traversing pre-operative scans (such as MRI or CT) can generate structurally accurate simulations as well as ground truth camera poses and depth maps. Although image-to-image (I2I) translation models such as CycleGAN can imitate realistic endoscopic images from these simulations, they are unsuitable for video-to-video synthesis due to the lack of temporal consistency, resulting in artifacts between frames. We propose MeshBrush, a neural mesh stylization method to synthesize temporally consistent videos with differentiable rendering. MeshBrush uses the underlying geometry of patient imaging data while leveraging existing I2I methods. With learned per-vertex textures, the stylized mesh guarantees consistency while producing high-fidelity outputs. We demonstrate that mesh stylization is a promising approach for creating realistic simulations for downstream tasks such as training networks and preoperative planning. Although our method is tested and designed for ureteroscopy, its components are transferable to general endoscopic and laparoscopic procedures. The code will be made public on GitHub [^1] https://github.com/juseonghan/MeshBrush.
@@
Mesh Stylization Endoscopy Differentiable Rendering 3D Style Transfer.
@@ 1 Introduction
Figure 1: Sample camera trajectory inside the mesh. From top to bottom: original rendered, I2I style transfers [19], and MeshBrush. Note inconsistency from frame to frame in the I2I method while MeshBrush maintains consistent features.
@@
To the best of our knowledge, our work is the first in 3D endoscopic mesh stylization to synthesize consistently style-transferred videos from rendered data. Because mesh textures are stationary, MeshBrush guarantees consistency regardless of time and view. Leveraging existing I2I style transfer modules, our self-supervised network simply requires an input mesh of the patient anatomy and a pretrained style transfer module. Our lightweight model produces high-resolution spatial textures on the patient mesh to create realistic endoscopic sequences with temporal and global consistency using a view-dependent heatmap loss. Finally, we demonstrate that temporally consistent style transfer is effective in vision algorithms such as feature matching and Structure-from-Motion.
@@ 2 Related Work
Few works address temporal and long-term consistent style transfer in medical scenes. Frisch et al. use optical flow (RAFT [27]) to warp views and impose a time-invariant structural similarity loss in cataract surgical video [8]. However, as flow estimates are imperfect, the results are suboptimal. Engelhardt et al. extend CycleGAN [31] to process multiple consecutive frames simultaneously for multi-frame consistency in phantom surgical videos [6], which cannot enforce long-term consistency, e.g. when the endoscope returns to a location after a certain duration. Rivoir et al. use the underlying 3D model to impose similar textures via a view consistency loss and learnable neural textures to ensure global consistency in simulated surgical scenes [21]. Although this method demonstrates remarkable performance, its applications are solely for laparoscopy which typically has simpler geometric structures than endoscopic anatomies. Mesh stylization ensures consistency regardless of time and viewpoint for complex 3D geometries such as simulated endoscopic environments.
@@
3 Methods
Preprocessing and Data Preparation. Fig. 2 shows the overview of MeshBrush. First, we segment a mesh from a CT scan with a medical visualization toolkit [^2] https://www.slicer.org/ and increase the vertex resolution by subdividing each triangle. This mesh is the structure upon which we will impose color information. We use an existing I2I method on specific camera renders and project these style transfer features onto the mesh. To generate meaningful mesh renders, we extract a mesh skeleton via [2], from which we generate $M$ camera positions with the farthest point down sampling. We set the camera view direction along its line segment of the skeleton to determine the $SE(3)$ camera poses $\{T_{i}\}_{i=1}^{M}$.
@@ 3 Methods
We generate supervisory I2I renderings as data preparation. From each camera pose $T_{i}$, we render a synthetic, unrealistic image $A_{i}$. Using a pre-trained image-to-image style transfer module [19], we translate $A_{i}$ into a realistic endoscopic view $B_{i}$. Our modularized approach allows our framework to be used for various medical domains as users can change the I2I style transfer method to match each domain. For each rendered view, we also compute a heatmap $H_{i}$, explained in the next section, which encodes the 3D scene geometry relative to the camera. We compile a set of $(T_{i},B_{i},H_{i})$ pairs for all $M$ views to use as ground truth supervision during the stylization process.
@@
Training. During training, our goal is to predict realistic endoscopic color values for each vertex of the mesh. We use a multilayer perception (MLP) network to regress RGB values; however, MLPs are not proficient in learning high-frequency associations such as texture variations. To overcome this issue, we first encode the vertex coordinates $V\in\mathbb{R}^{N\times 3}$ of the mesh to a Fourier feature encoding [26], i.e. $\gamma(V)=(\cos(2\pi GV),\sin(2\pi GV))$, where $G$ is a matrix with elements drawn from a zero-mean Gaussian distribution. The encoded vertices $\gamma(V)$ are passed into the MLP to generate texture predictions, which are applied onto the mesh. We then re-render a new set of images $\hat{A}_{i}$ from camera views $T_{i}$ from the data preparation phase. We supervise $\hat{A}_{i}$ with its corresponding style transferred image $B_{i}$ via the view-dependent loss, expounded in the next section. The MLP receives gradient updates from the error of $\hat{A}_{i}$ and $B_{i}$ via differentiable rendering. We repeat this over all $M$ camera views to complete one epoch of optimization. Over multiple epochs, the error between $\hat{A}_{i}$ and $B_{i}$ is minimized through training, which causes the mesh textures to mimic the style-transferred images.
@@
Using a neural network to predict textures rather than projecting the style transfer RGB values to the mesh has two important advantages. First, vertices are much spatially finer than pixels, which enables the mesh to learn more realistic textures at a higher resolution. Second, projecting RGB values from $B_{i}$ introduces quilting, where camera poses with overlapping image regions introduce sharp texture discontinuities due to inconsistency in the transferred styles.
We omit the subscript $i$ in this section for the sake of clarity. Given a rendered image $A$, we calculate a normal map which contains the vertex normals at each pixel of the rendered image. By computing the dot product between the normal map and the camera viewing direction, we can calculate the cosine of the angle between each pixel’s 3D point and the camera orientation, creating the view-orientation heatmap $\theta$. Similarily, the view-distance heatmap $D$ can be calculated by the $L_{2}$ distance between the camera position and 3D points in the rendered image. After scaling $D$ to unit, the heatmaps are aggregated to create the final heatmap $H\in\left[0,1\right]$.
$H=\frac{1}{2}\left(\frac{1-\theta}{2}+D\right)$ (1)
The final heatmap is generated by forcing pixels with view-direction values greater than a user-defined threshold to 0, and $A$ and $B$ are converted into LAB color space for more accurate color predictions. This results in our final heatmap-weighed mean squared error loss function:
$\mathcal{L}=H\cdot|A-B|^{2}_{2}$ (2)
where $A$ is a rendered image from the mesh and $B$ is its corresponding style transferred image acquired during the data preparation phase.
@@
4 Experiments and Results
@@ 5 Discussion and Conclusion
Structure-from-Motion. The sparse reconstructions of COLMAP, for both PyTorch3D and Pyrender renderings of the mesh, failed for both untextured and I2I methods. This confirms that textures from untextured and I2I-translated images do not provide unique descriptors for global camera localization. On the other hand, the features predicted by MeshBrush have globally distinct descriptors, shown in Fig. 4. We found that COLMAP uses 41.3% from 1,120 images to reconstruct 32327 points. Previous work achieved an average of 20.6% from 498 images to reconstruct 4838 points for kidney reconstructions from rendered views [1]. Thus, our method ensures the global and long-term consistency necessary for sophisticated downstream vision tasks like 3D reconstruction.
@@
{credits}
5.0.1 Acknowledgements
This study was partially supported by the NIH T32 Training Grant (grant number T32EB021937) and the Vanderbilt Institute of Surgery and Engineering Seed Grant.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.